How many kilograms of milk testing 5% butterfat and how many kilograms of cream testing 25% butterfat should be mixed to give 100 kilograms testing 20% butterfat ? Not only show me how it's done , but tell me why the certain steps are taken . Thanks alot !!!
Answer by guapa(62)

Let x represent the amount of 5% milk butterfat. Then 100-x represents the amount of 25% cream butterfat.
Use the following guideline:
Kg of 5% milk butterfat + Kg of 25% cream butterfat = Kg of 20% mix
5%(x) + 25%(100-x) = 20%(100)
Now let's solve this equation:
0,05(x)+ 0,25(100-x)= 0,2(100) Now simplify the terms
0,05x + 25 - 0,25x = 20 combine like terms
-0,2x +25 = 20 Substract 25 from both sides
-0,2x +25 -25 = 20-25
-0,2x = -5 Divide both sides by -0,2
-0,2x/-0,2 = -5/-0,2
x = 25
Therefore, 100-x = 100-25 = 75
We should mix 25kg of milk butterfat with 75kg of cream butterfat.
I hope this helps